# IUST_PersonReId: A New Domain in Person Re-Identification Datasets

### v1

@@ 4.6 Ablation Studies
This setup examines how well models differentiate gender-based features. Initially, the models were trained on the original dataset, which is imbalanced, with a majority of male identities (1,338) compared to female identities (509). To investigate the effect of this imbalance, we conducted an additional experiment where the dataset was balanced by undersampling male identities to match the number of female identities. The models were then fine-tuned on the balanced dataset. Results from both the original and balanced datasets are presented in Tables 7 and 8, respectively.
Table 7: Gender Comparison on Original Dataset
Model Gender Rank-1 Rank-5 Rank-10 mAP
Solider [31] Male 59.79% 73.03% 78.05% 49.2%
CLIP-ReID [32] 68.3% 78.9% 82.8% 62.21%
Solider [31] Female 34.98% 51.08% 57.27% 29.67%
CLIP-ReID [32] 45.6% 57.9% 65.3% 36.27%
Table 8: Gender Comparison on Balanced Dataset
Model Gender Rank-1 Rank-5 Rank-10 mAP
Solider [31] Male 40.44% 43.56% 48.00% 38.42%
CLIP-ReID [32] 69.57% 82.34% 84.78% 64.45%
Solider [31] Female 30.34% 37.15% 40.87% 24.98%
CLIP-ReID [32] 44.68% 58.80% 65.74% 35.01%
@@
    Clear: Images where all face and upper body keypoints are visible, with at most 4 keypoints missing.
@@
    Partial: Images where at least half of the 18 keypoints are visible.
@@
    Occluded: Images where fewer than half of the 18 keypoints are visible.
Impact of Visibility on Re-identification Performance
These categories were determined based on the visibility and clarity of keypoints, reflecting variations in occlusion, pose, or camera angle. Performance metrics were computed separately for each visibility category to evaluate the models’ ability to generalize across varying levels of complexity. Results of this study are presented in Table 9.
Table 9: Visibility Comparison
Model Visibility Rank-1 Rank-5 Rank-10 mAP
Solider [31] Clear 56.11% 70.89% 76.56% 45.85%
CLIP-ReID [32] 66.67% 77.62% 81.62% 58.84%
Solider [31] Partial 55.66% 69.81% 71.69% 48.12%
CLIP-ReID [32] 65.45% 72.12% 79.39% 60.25%
Solider [31] Occluded 24.30% 36.11% 43.05% 21.80%
CLIP-ReID [32] 32.39% 47.89% 55.87% 27.75%
Based on the results, we can conclude that the model’s performance is highly sensitive to the visibility of the individual’s body. When more keypoints are visible, the re-identification task becomes significantly easier. In contrast, when keypoints are less visible due to occlusion or challenging poses, the re-identification becomes much more difficult, demonstrating the critical role of visibility in person re-identification tasks.

### v2

@@ 4.4.2 Gender-based Query and Gallery Setup
This setup examines how well models differentiate gender-based features. Initially, the models were trained on the original dataset, which is imbalanced, with a majority of male identities (1,338) compared to female identities (509). To investigate the effect of this imbalance, we conducted an additional experiment where the dataset was balanced by undersampling male identities to match the number of female identities. The models were then fine-tuned on the balanced dataset. Results from both the original and balanced datasets are presented in Table 7.
Table 7: Gender-wise Performance Comparison Between Original and Balanced Datasets for Each Model
Gender Model Rank-1 Rank-5 Rank-10 mAP
Orig. Bal. Orig. Bal. Orig. Bal. Orig. Bal.
Male SOLIDER [31] 59.79% 40.44% 73.03% 43.56% 78.05% 48.00% 49.20% 38.42%
CLIP-ReID [32] 68.30% 69.57% 78.90% 82.34% 82.80% 84.78% 62.21% 64.45%
Female SOLIDER [31] 34.98% 30.34% 51.08% 37.15% 57.27% 40.87% 29.67% 24.98%
CLIP-ReID [32] 45.60% 44.68% 57.90% 58.80% 65.30% 65.74% 36.27% 35.01%
@@ 4.4.3 Impact of Visibility on Re-identification Performance
    Clear: Most of the face and upper body are visible, with no more than 4 missing keypoints.
@@
    Partial: At least 9 out of 18 keypoints are visible.
@@
    Occluded: Fewer than 9 out of 18 keypoints are visible.
These categories were determined based on the visibility and clarity of keypoints, reflecting variations in occlusion, pose, or camera angle. Mean average precision was computed separately for each visibility category and gender to evaluate the model’s performance under varying visual conditions. As shown in Figure 5, the model achieves the highest mAP on clear examples, with a noticeable performance drop under partial and occluded settings.
Figure 5: Mean Average Precision of CLIP-ReID Across Visibility Categories by Gender.
Table 8: Number of Male and Female Queries per Visibility Category.
Category #Male #Female
Clear 401 83
Partial 58 21
Occluded 17 74
Interestingly, the gender-wise breakdown reveals a performance disparity linked to the distribution of identities across visibility categories. Table 8 shows that clear queries are predominantly male, who are generally more visible in the dataset. In contrast, the majority of occluded queries are female, often due to cultural attire such as the hijab, which contributes to increased occlusion and reduced keypoint visibility. This skew in data distribution helps explain why female performance is lower under the clear setting but surpasses male performance under partial visibility, where fewer clear cues are available for either gender. Overall, these findings highlight the critical role of visibility and gender-specific occlusion in the re-identification task.
4.4.4 Facial Feature Analysis
To investigate the role of facial features in person re-identification, we conducted a series of experiments focused on face visibility and its potential integration with full-body cues. First, we examined the impact of face blurring, which was applied to the publicly released version of our dataset for privacy and ethical considerations. As shown in Figure 7, the results indicate that blurring faces leads to only a marginal reduction in performance, suggesting that facial features, while informative in certain cases, are not the dominant cues leveraged by person ReID models in real-world surveillance scenarios.
Next, we created a subset of the test set where the face was clearly visible in the query image and at least 10% of the corresponding gallery images. We performed three separate experiments on this subset: (1) using Additive Angular Margin Loss (ArcFace [36]), a face recognition model to extract facial embeddings and perform matching based on Euclidean distance, (2) using a full-body person re-identification model (CLIP-ReID) with the same matching protocol, and (3) applying a fusion strategy that combined the distance matrices from both models to jointly leverage facial and body cues. As illustrated in Figure 7, the person re-identification model outperformed the face-only approach, and the fusion of face and body distances slightly degraded the overall accuracy. This degradation is attributed to the limitations inherent in surveillance footage, such as low face resolution, suboptimal lighting, and non-frontal angles, which reduce the reliability of facial features.
Figure 6: Impact of face blurring on the person re-identification performance.
Figure 7: Evaluation of face recognition, person re-identification, and their combined fusion performance.
These findings emphasize that in realistic, in-the-wild settings, especially those involving modest attire and partially occluded faces robust person re-identification is best achieved through comprehensive body-based feature representations rather than relying heavily on facial cues.